Problem 21.83 Part A Packing material made of pieces of foamed polystyrene can easily become charged and stick to each other. Given that the density of this material is about 35 kg/m1, estimate how much charge might be on a 2.0 cm-diameter foamed polystyrene sphere, assuming the electric force between two spheres stuck together is equal to the weight of one sphere Express your answer using two significant figures Submit My Answers Give Up Provide Feedback Continue
Answer #1

Weight of each sphere , W = density*volume*g

= 35*((4/3)*pi*0.01^3)*9.8

= 0.00144 N

So, electric force between them, Fe = 0.00144 N

Now, Fe = k*Q^2/d^2

where: d = 2 cm = 0.02 m

k = 9*10^9

So, 9*10^9*Q^2/0.02^2 = 0.00144

So, Q = 8*10^-9 C <--------answer